# Quantum two-dimensional superintegrable systems in flat space: exact-solvability, hidden algebra, polynomial algebra of integrals

### v1

@@ Abstract
In this short review paper the detailed analysis of six two-dimensional quantum superintegrable systems in flat space is presented. It includes the Smorodinsky-Winternitz potentials I-II (the Holt potential), the Fokas-Lagerstrom model, the 3-body Calogero and Wolfes (equivalently, $G_{2}$ rational, or $I_{6}$) models, and the Tremblay-Turbiner-Winternitz (TTW) system with integer index $k$. It is shown that all of them are exactly-solvable, thus, confirming the Montreal conjecture (2001); they admit algebraic forms for the Hamiltonian and both integrals (all three can be written as differential operators with polynomial coefficients without a constant term), they have polynomial eigenfunctions with the invariants of the discrete symmetry group of invariance taken as variables, they have hidden (Lie) algebraic structure $g^{(k)}$ with various $k$, and they possess a (finite order) polynomial algebras of integrals. Each model is characterized by infinitely-many finite-dimensional invariant subspaces, which form the infinite flag. Each subspace coincides with the finite-dimensional representation space of the algebra $g^{(k)}$ for a certain $k$. In all presented cases the algebra of integrals is a 4-generated $(H,I_{1},I_{2},I_{12}\equiv[I_{1},I_{2}])$ infinite-dimensional algebra of ordered monomials of degrees 2,3,4,5, which is a subalgebra of the universal enveloping algebra of the hidden algebra.
@@ Introduction
This review paper is organized as follows. In Section I the general theory of (maximally) superintegrable systems is presented. Section II is about the 4-generated polynomial algebra of integrals and its representation theory in differential operators for planar superintegrable systems. The algebra $g^{(s)}$ is described in Section III. Various 2D superintegrable systems listed in Bonatsos et al, [6, 7] are contained in Section IV: Smorodinsky-Winternitz Case I (Part A) and Case II (Part B), Fokas-Lagerstrom (Part C). In Section V the celebrated 3-body Calogero model, equivalently, ${A}_{2}$ rational model at $\omega=0$ (singular case) and in Section VI the 3-body Wolfes model, equivalently, ${G}_{2}/I_{6}$ rational model at $\omega=0$ (singular case). Section VII is about the TTW system with integer index $k$. Section VIII is written by AVT and dedicated to the memory of Pavel Winternitz, the 3rd coauthor who recently passed away.
@@ I Generalities
$g^{\mu\nu}(q)\,p_{\mu}\,p_{\nu}\ \rightarrow\ \sqrt{g}\,\partial_{\mu}\,g^{\mu\nu}\,\frac{1}{\sqrt{g}}\,\partial_{\nu}\ ,$
@@
$\Delta_{g}\ =\ \sqrt{g}\,\partial_{\mu}\,g^{\mu\nu}\,\frac{1}{\sqrt{g}}\,\partial_{\nu}\ ,$ (9)
where $g=\det g^{\mu\nu}$, $\partial_{\mu}=\frac{\partial}{\partial{x_{\mu}}}$ and $V$ is a potential function on the manifold. We are mostly focussed on flat space, where the Riemann tensor for $g_{\mu\nu}$ vanishes. For two-dimensional space this corresponds to zero curvature.
@@
  • The quantum completely-integrable system is superintegrable if there are $s\geq 1$ additional partial differential operators with variable coefficients ${\cal I}_{n},{\cal I}_{n+1},\cdots,{\cal I}_{n+s-1}$ such that ${\cal H},{\cal I}_{j}]=0$ for $0\leq j\leq n+s-1$ and the set $\{{\cal I}_{j}:\ 0\leq j\leq n+s-1\}$ is algebraically independent. If s=n1s=n-1, apparently the maximum possible, the system is maximally superintegrable. If the additional integrals are chosen to be the differential operators of a minimal order, they are called basic additional quantum integrals. In general, additional integrals do not commute with each other or with integrals from the commutative algebra ${\cal I}_{i}$ for $1\leq i\leq n-1$. The basic integrals ${\cal I}_{j}$ of a superintegrable system can generate a polynomial algebra under the Lie bracket/commutator, which is not usually a Lie algebra. However, if all basic integrals are first order differential operators, they span a Lie algebra. The simplest example of this situation is the isotropic harmonic oscillator.
@@ IV.1 Smorodinsky-Winternitz system, Case I (1965)
$i_{12}\ \equiv\ \Gamma^{-1}\,{\cal I}_{12}\,\Gamma$
@@
see [4]. All four operators (31)-(IV.1) have no constant terms (the terms containing no derivatives). Their lowest eigenfunctions are constants and the lowest eigenvalues are zeroes. Evidently, $i_{1},i_{2},i_{12}$ commute with the algebraic Hamiltonian $h$, their double commutators are at most of the 2nd degree polynomials in $h,i_{1},i_{2},i_{12}$ but without constant terms, cf. (27)-(28). All four algebraic operators $h,i_{1},i_{2},i_{12}$ are elements of the universal enveloping algebra of the 5-dimensional Heisenberg algebra in the coordinate-momentum representation $(\partial_{\tau},\tau)$.
@@ IV.2 Smorodinsky-Winternitz system, Case II (1965), Holt model (1982)
$h\ \equiv\ \Gamma^{-1}\,(H\,-\,E^{(k)}_{0})\,\Gamma$
$\ =\ -2\,\tau_{1}\,\partial^{2}_{\tau_{1}}\ -\ \frac{1}{2}\,\partial^{2}_{y}\ -\ (\beta_{k}\,-\,2\,\sqrt{2}\,\omega\,\tau_{1})\,\partial_{\tau_{1}}\ +\ 2\,\sqrt{2}\,\omega\,y\,\partial_{y}\ ,$ (48)
@@
see [4]. It implies that the SW-II system has the hidden algebra $sl(3,R)\in g^{(1)}$ similar to SW-I. Spectra of $h$ (IV.2):
@@
cf.(47), which are invariants of the reflection group. (56) is the ground state eigenfunction of (40) written in coordinates $\tau_{1},\tau_{2}$. Then it has to be made the gauge rotation. As for the Hamiltonian in algebraic form it is, in fact, (IV.2) with $y$ replaced by $\tau_{2}=y^{2}$,
@@ IV.4 Fokas-Lagerstrom potential (1980)
$i_{2}\ \equiv\ \Gamma^{-1}\,{\cal I}_{2}\,\Gamma$ (71)
@@
However, in order to get the second integral in the algebraic form we have to square the integral $i_{2}$ (71) and then change variables $(x,y)$ to $\tau_{1,2}$:
@@ VIII History: to the memory of Pavel Winternitz (AVT)
When a monumental review paper by Miller-Post-Winternitz [3] was published, Pavel Winternitz approached one of us (AVT) saying that after a critical analysis of this review [3] it becomes clear that (I) the formalism of quantum superintegrability should be further elaborated emphasizing the polynomial algebras of integrals and (II) the most important examples should be profoundly studied, in particular, the so-called Smorodinsky-Winternitz potentials I-II [1]. During 2013-2016 Pavel Winternitz and AVT worked out the formalism (I) and by 2018 the Introduction and Generalities were mainly written in the present form  [^12] We preferred to make minimal modifications of these two parts as a memory to P Winternitz. and we proceeded (slowly, without rush) to the Smorodinsky-Winternitz potentials I-II. After the sudden death of P Winternitz in 2021 the review was continued with JC Lopez Vieyra and eventually completed by him together with AVT.
@@ Conclusions
The analysis of three quantum superintegrable systems in non-flat space (the Higgs I-II potentials, thus, the spherical/hyperbolic harmonic oscillator and the Kepler-Coulomb problem [24], the reduced Coulomb problem [35]) supports the validity of the original Montreal conjecture - they are exactly-solvable with explicitly known energy eigenvalues, they also possess the polynomial algebras of integrals [6], [35].
@@ Acknowledgments
A.V.T. is thankful to CRM, University of Montreal for its kind hospitality extended to him where this work was started and progressed in 2013-6 together with P Winternitz, who passed away on Febr.13, 2021, and it never was completed. The authors discussed the subject with W Miller Jr (University of Minnesota) on multiple occasions, who passed away on Oct.29, 2023. This paper is dedicated to the memory of Pavel Winternitz and Willard Miller Jr, two giants of mathematical physics. Both present authors feel strongly their absence.
AVT thanks G Sterman, R Schrock, B McCoy (YITP, Stony Brook) for their kind invitation to present a regular seminar, where the idea to write a review about algebras of integrals was born - it is a part of the present review. The authors thank Dr K K Phua (World Scientific) for the kind invitation to write this review paper. We are thankful to Dr A M Escobar-Ruiz for participation in the early stage of work and numerous discussions. JCLV was supported in part by PASPA project for his sabbatical stay in Southern Methodist University (Dallas, USA), where a part of this review was written.
This work is supported in part by the the PAPIIT grants IN108815, IN113022, IN104125 and CONACyT grants 166189, A1-S-17364 (Mexico).

### v2

@@ Abstract
In this short review paper the detailed analysis of six two-dimensional quantum superintegrable systems in flat space is presented. It includes the Smorodinsky-Winternitz potentials I-II (the Holt potential), the Fokas-Lagerstrom model, the 3-body Calogero and Wolfes (equivalently, $G_{2}$ rational, or $I_{6}$) models, and the Tremblay-Turbiner-Winternitz (TTW) system with integer index $k$. It is shown that all of them are exactly-solvable, thus, confirming the Montreal conjecture (2001); they admit algebraic forms for the Hamiltonian and both integrals (all three can be written as differential operators with polynomial coefficients without a constant term), they have polynomial eigenfunctions with the invariants of the discrete symmetry group of invariance taken as variables, they have hidden (Lie) algebraic structure $g^{(k)}$ with various $k$, and they possess a (finite order) polynomial algebras of integrals. Each model is characterized by infinitely-many finite-dimensional invariant subspaces, which form the infinite flag. Each subspace coincides with the finite-dimensional representation space of the algebra $g^{(k)}$ for a certain $k$. In all presented cases the algebra of integrals is a 4-generated $(H,{\cal I}_{1},{\cal I}_{2},{\cal I}_{12}\equiv[{\cal I}_{1},{\cal I}_{2}])$ infinite-dimensional algebra of ordered monomials of degrees 2,3,4,5, which is a subalgebra of the universal enveloping algebra of the hidden algebra.
@@ Introduction
This review paper is organized as follows. In Section I the general theory of (maximally) superintegrable systems is presented. Section II is about the 4-generated polynomial algebra of integrals and its representation theory in differential operators for planar superintegrable systems. The algebra $g^{(s)}$ is described in Section III. Various 2D superintegrable systems listed in Bonatsos et al, [6, 7] are contained in Section IV: Smorodinsky-Winternitz Case I (Part A) and Case II (Part B), Fokas-Lagerstrom (Part C). In Section V the celebrated 3-body Calogero model, equivalently, ${A}_{2}$ rational model at $\omega=0$ (singular case) and in Section VI the 3-body Wolfes model, equivalently, ${G}_{2}/I_{6}$ rational model at $\omega=0$ (singular case). Section VII is about the TTW system with integer index $k$. Section VIII is written by AVT and dedicated to the memory of Pavel Winternitz, the 3rd coauthor, who recently passed away.
@@ I Generalities
$g^{\mu\nu}(q)\,p_{\mu}\,p_{\nu}\ \rightarrow\ \frac{1}{\sqrt{g}}\,\partial_{\mu}\,g^{\mu\nu}\,{\sqrt{g}}\,\partial_{\nu}\ ,$
@@
$\Delta_{g}\ =\ \frac{1}{\sqrt{g}}\,\partial_{\mu}\,g^{\mu\nu}\,{\sqrt{g}}\,\partial_{\nu}\ ,$ (9)
where $g=\det g_{\mu\nu}$, $\partial_{\mu}=\frac{\partial}{\partial{x_{\mu}}}$ and $V$ is a potential function on the manifold. We are mostly focussed on flat space, where the Riemann tensor for $g_{\mu\nu}$ vanishes. For two-dimensional space this corresponds to zero curvature.
@@
  • The quantum completely-integrable system is superintegrable if there are $s\geq 1$ additional partial differential operators with variable coefficients ${\cal I}_{n},{\cal I}_{n+1},\cdots,{\cal I}_{n+s-1}$ such that ${\cal H},{\cal I}_{j}]=0$ for $0\leq j\leq n+s-1$ and the set $\{{\cal I}_{j}:\ 0\leq j\leq n+s-1\}$ is algebraically independent. If s=n1s=n-1, apparently the maximum possible, the system is maximally superintegrable. If the additional integrals are chosen to be the differential operators of a minimal order, they are called basic additional quantum integrals. In general, additional integrals do not commute with each other, ${\cal I}_{n+i},{\cal I}_{n+j}]\neq 0$ for $i\neq j\geq 0$, or with integrals from the commutative algebra ${\cal I}_{i}$ for $1\leq i\leq n-1$, ${\cal I}_{i},{\cal I}_{n+j}]\neq 0$ for $j\geq 0$. In particular, for two-dimensional case ${\cal I}_{1},{\cal I}_{2}]\equiv{\cal I}_{12}\neq 0$. The basic integrals ${\cal I}_{j}$ of a superintegrable system can generate a polynomial algebra under the Lie bracket/commutator, which is not usually a Lie algebra. However, if all basic integrals are first order differential operators, they span a Lie algebra. The simplest example of this situation is the isotropic harmonic oscillator.
@@ IV.1 Smorodinsky-Winternitz system, Case I (1965)
$i_{12}\ \equiv\ \Gamma^{-1}\,{\cal I}_{12}\,\Gamma$ (34)
@@
see [4]. All four operators (31)-(34) have no constant terms (the terms containing no derivatives). Their lowest eigenfunctions are constants and the lowest eigenvalues are zeroes. Evidently, $i_{1},i_{2},i_{12}$ commute with the algebraic Hamiltonian $h$, their double commutators are at most of the 2nd degree polynomials in $h,i_{1},i_{2},i_{12}$ but without constant terms, cf. (27)-(28). All four algebraic operators $h,i_{1},i_{2},i_{12}$ are elements of the universal enveloping algebra of the 5-dimensional Heisenberg algebra in the coordinate-momentum representation $(\partial_{\tau},\tau)$.
@@ IV.2 Smorodinsky-Winternitz system, Case II (1965), Holt model (1982)
$h\ \equiv\ \Gamma^{-1}\,(H\,-\,E^{(k)}_{0})\,\Gamma$ (48)
$\ =\ -2\,\tau_{1}\,\partial^{2}_{\tau_{1}}\ -\ \frac{1}{2}\,\partial^{2}_{y}\ -\ (\beta_{k}\,-\,2\,\sqrt{2}\,\omega\,\tau_{1})\,\partial_{\tau_{1}}\ +\ 2\,\sqrt{2}\,\omega\,y\,\partial_{y}\ ,$
@@
see [4]. It implies that the SW-II system has the hidden algebra $sl(3,R)\in g^{(1)}$ similar to SW-I. Spectra of $h$ (48):
@@
cf.(47), which are invariants of the reflection group. (56) is the ground state eigenfunction of (40) written in coordinates $\tau_{1},\tau_{2}$. Then it has to be made the gauge rotation. As for the Hamiltonian in algebraic form it is, in fact, (48) with $y$ replaced by $\tau_{2}=y^{2}$,
@@ IV.4 Fokas-Lagerstrom potential (1980)
$i_{2}\ \equiv\ \Gamma^{-1}\,{\cal I}_{2}\,\Gamma$
@@
However, in order to get the second integral in the algebraic form we have to square the integral $i_{2}$ (IV.4) and then change variables $(x,y)$ to $\tau_{1,2}$:
@@ VIII History: to the memory of Pavel Winternitz (AVT)
When a monumental review paper by Miller-Post-Winternitz [3] was published, Pavel Winternitz approached one of us (AVT) saying that after a critical analysis of this review [3] it becomes clear that (I) the formalism of quantum superintegrability should be further elaborated emphasizing the polynomial algebras of integrals and (II) the most important examples should be profoundly studied, in particular, the so-called Smorodinsky-Winternitz potentials I-II [1]. During 2013-2016 Pavel Winternitz and AVT worked out the formalism (I) and by 2018 the Introduction and Generalities were mainly written in the present form   [^12] We preferred to make minimal modifications of these two parts as a memory to P Winternitz. and we proceeded (slowly, without rush) to the Smorodinsky-Winternitz potentials I-II. After the sudden death of Pavel Winternitz in 2021 the review was continued with JC Lopez Vieyra and eventually completed by him together with AVT.
@@ Conclusions
It must be noted that the existence of the algebraic form of superintegrable system allows us to construct the Fock space representation in the operators/letters $a_{1,2},b_{1,2}$, which are elements of five-dimensional Heisenberg algebra $h_{5}$. These operators can be realized as finite-difference and/or shift operators on uniform or exponential lattice. Thus, we arrive at the isospectral systems on uniform, exponential lattices and mixed, uniform-exponential lattices, see e.g. [36] and references therein.
The analysis of three quantum superintegrable systems in non-flat space (the Higgs I-II potentials, thus, the spherical/hyperbolic harmonic oscillator and the Kepler-Coulomb problem [24], the reduced Coulomb problem [35]) supports the validity of the original Montreal conjecture - they are exactly-solvable with explicitly known energy eigenvalues, they also possess the polynomial algebras of integrals [6], [35]. Recently, a new generalization of the TTW system into the spaces of constant curvature $S_{2}$ and $H_{2}$ was proposed [37]: it remained superintegrable (both integrals might be constructed explicitly at least for some index $k$) and exactly-solvable, while the question about polynomiality of the algebra of integrals remained open.
@@ Acknowledgments
AVT is thankful to CRM, University of Montreal for its kind hospitality extended to him where this work was started and progressed in 2013-6 together with P Winternitz, who passed away on Febr.13, 2021, and it never was completed. The authors discussed the subject with W Miller Jr (University of Minnesota) on multiple occasions, who passed away on Oct.29, 2023. This paper is dedicated to the memory of Pavel Winternitz and Willard Miller Jr, two giants of mathematical physics. Both present authors feel strongly their absence.
AVT expresses the gratitude to G Sterman, R Schrock, B McCoy (YITP, Stony Brook) for their interest to the subject and kind invitation to present a regular seminar at YITP on this subject, where the idea to write a review about algebras of integrals was born - it is a part of the present review. The authors thank Dr K K Phua (World Scientific) for the kind invitation to write this review paper. We are thankful to A M Escobar-Ruiz for participation in the early stage of work and numerous discussions. JCLV was supported in part by PASPA project for his sabbatical stay in Southern Methodist University (Dallas, USA), where significant part of this review was prepared.
This work is supported in part by the PAPIIT grants IN108815, IN113022, IN104125 and CONACyT grants 166189, A1-S-17364 (Mexico).